# Towards Scaling Laws for Symbolic Regression

### v1

@@ 3 Methodology
Data generation Fig. 1 shows our two-step approach for data generation. While previous work usually adopts the expression sampling mechanism proposed by Lample and Charton [9], we generate a complete set of expressions recursively, starting with just the variables and then iteratively applying unary and binary operators. This allows us to get more control over the expressions the model sees during training, especially in terms of the number of duplicates, and avoids biasing the model towards certain, more common expressions. Furthermore, we filter out equivalent expressions and unify their representation, allowing us to feed the model with cleaner training data. We obtain our set of base expressions $E$ by taking all expressions up to a fixed threshold. For each of these expressions, we try to sample $k$ expression-dataset pairs. Therefore, we first insert random constants and then sample a random dataset from a Gaussian mixture following Kamienny et al. [2]. More details about our data generation can be found in App. A.
@@ 4.1 Setup
We found that the AdamCPR optimizer [15], combined with a linear warm-up of the learning rate for the first 5% of steps and cosine annealing thereafter, works best for our setting. All training hyperparameters are reported in App. C. For evaluation, we first randomly sample $128$ expressions from the model and keep the one with the highest $R^{2}$ (coefficient of determination) score. Then, we look at both the perfect-solved ratio, denoted as $\text{Acc}_{\text{solved}}$, and the expression ratio with $R^{2}>0.99$, denoted as $\text{Acc}_{R^{2}>0.99}$, on the test split. We run each evaluation over three different seeds and report the mean $\text{Acc}_{\text{solved}}$ and $\text{Acc}_{R^{2}>0.99}$ across these seeds.
@@ 4.2 Results
We report the results of our hyperparameter sweep in App. B (Fig. 4). Detailed evaluation results of the trainings with the best found hyperparameters can be found in App. B (Fig. 2). In the following, we highlight our three key findings.
@@
Solved rate and loss scale as power laws with compute To analyze the effect of scaling on performance, we look at training with the optimal configuration per model size. Following the approach of Hoffmann et al. [6], we get a Pareto front of our models by binning compute into $1,500$ intervals and keeping the models with the lowest validation loss up to their compute level. We adopt the assumption that scaling trends follow power-law trends and thus fit power laws on the remaining data. Fig. 3 shows the results for the $\text{Acc}_{\text{solved}}$ metric, which follows power-law scaling with compute across three orders of magnitude. In App. B, we report analogous plots for $\text{Acc}_{R^{2}>0.99}$ and test loss, yielding similar insights. Larger models with more data consistently improve the solved ratio and reduce loss, with no signs of saturation in the largest models tested. For instance, we can see that $\text{Acc}_{\text{solved}}$ increases from approximately $0.03$ at the lowest compute budget to $0.6$ at the highest compute budget, and that according to the scaling law, we could reach $0.8$ with a compute budget of $3.8\times 10^{21}$ FLOPs. Also, by comparing the scaling laws, we can see that improvements in $\text{Acc}_{R^{2}>0.99}$ happen much faster than in $\text{Acc}_{\text{solved}}$, highlighting how hard it is to achieve exact expression matches.
Optimal batch size and learning rate grow with compute Based on the results of our grid search, we analyze the scaling trends of the optimal learning rate and batch size. We follow the strategy of Porian et al. [14], which involves two-step interpolation by first interpolating the optimal learning rates and then the optimal batch sizes using Akima interpolation, followed by fitting scaling laws to both the optimal interpolated values. App. B (Fig. 5) shows systematic trends for the compute-optimal choices of batch size and learning rate, indicating both should increase when scaling compute. The upward trend for the learning rate contrasts with findings for LLMs, showcasing how training characteristics can vary between different tasks. While we can see clear tendencies, we also note that more training runs over multiple seeds are needed to further reduce variance.
Optimal trade-off between model size and data size We fit scaling laws to the optimal number of parameters and expression tokens per compute budget. The results are shown in App. B (Fig. 6). We look at the ratio of the predicted scaling laws, noticing an increasing trend with growing compute. Although a token-to-parameter ratio of $\approx$15 seems to be optimal in our compute regime, the results indicate that the training dataset size should scale slightly faster than the model size.
@@ Acknowledgements
Appendix A Data generation details
A.1 Expression generation
@@ A.2 Data sampling
Appendix B Additional Results
@@
Appendix C Training Details

### v2

@@ 3 Methodology
Data generation Fig. 1 shows our two-step approach for data generation. While previous work usually adopts the expression sampling mechanism proposed by Lample and Charton [8], we generate a complete set of expressions recursively, starting with just the variables and then iteratively applying unary and binary operators. This allows us to get more control over the expressions the model sees during training, especially in terms of the number of duplicates, and avoids biasing the model towards certain, more common expressions. Furthermore, we filter out equivalent expressions and unify their representation, allowing us to feed the model with cleaner training data. We obtain our set of base expressions $E$ by taking all expressions up to a fixed threshold. For each of these expressions, we try to sample $k$ expression-dataset pairs. Therefore, we first insert random constants and then sample a random dataset from a Gaussian mixture following Kamienny et al. [6]. More details about our data generation can be found in App. B.
@@ 4.1 Setup
We found that the AdamCPR optimizer [2], combined with a linear warm-up of the learning rate for the first 5% of steps and cosine annealing thereafter, works best for our setting. All training hyperparameters are reported in App. D. For evaluation, we first randomly sample $128$ expressions from the model and keep the one with the highest $R^{2}$ (coefficient of determination) score. Then, we look at both the perfect-solved ratio, denoted as $\text{Acc}_{\text{solved}}$, and the expression ratio with $R^{2}>0.99$, denoted as $\text{Acc}_{R^{2}>0.99}$, on the test split. We run each evaluation over three different seeds and report the mean $\text{Acc}_{\text{solved}}$ and $\text{Acc}_{R^{2}>0.99}$ across these seeds.
@@ 4.2 Results
We report the results of our hyperparameter sweep in App. C (Fig. 4). Detailed evaluation results of the trainings with the best found hyperparameters can be found in App. C (Fig. 2). In the following, we highlight our three key findings.
@@
Solved rate and loss scale as power laws with compute To analyze the effect of scaling on performance, we look at training with the optimal configuration per model size. Following the approach of Hoffmann et al. [4], we get a Pareto front of our models by binning compute into $1,500$ intervals and keeping the models with the lowest validation loss up to their compute level. We adopt the assumption that scaling trends follow power-law trends and thus fit power laws on the remaining data. Fig. 3 shows the results for the $\text{Acc}_{\text{solved}}$ metric, which follows power-law scaling with compute across three orders of magnitude. In App. C, we report analogous plots for $\text{Acc}_{R^{2}>0.99}$ and test loss, yielding similar insights. Larger models with more data consistently improve the solved ratio and reduce loss, with no signs of saturation in the largest models tested. For instance, we can see that $\text{Acc}_{\text{solved}}$ increases from approximately $0.03$ at the lowest compute budget to $0.6$ at the highest compute budget, and that according to the scaling law, we could reach $0.8$ with a compute budget of $3.8\times 10^{21}$ FLOPs. Also, by comparing the scaling laws, we can see that improvements in $\text{Acc}_{R^{2}>0.99}$ happen much faster than in $\text{Acc}_{\text{solved}}$, highlighting how hard it is to achieve exact expression matches.
Optimal batch size and learning rate grow with compute Based on the results of our grid search, we analyze the scaling trends of the optimal learning rate and batch size. We follow the strategy of Porian et al. [13], which involves two-step interpolation by first interpolating the optimal learning rates and then the optimal batch sizes using Akima interpolation, followed by fitting scaling laws to both the optimal interpolated values. App. C (Fig. 5) shows systematic trends for the compute-optimal choices of batch size and learning rate, indicating both should increase when scaling compute. The upward trend for the learning rate contrasts with findings for LLMs, showcasing how training characteristics can vary between different tasks. While we can see clear tendencies, we also note that more training runs over multiple seeds are needed to further reduce variance.
Optimal trade-off between model size and data size We fit scaling laws to the optimal number of parameters and expression tokens per compute budget. The results are shown in App. C (Fig. 6). We look at the ratio of the predicted scaling laws, noticing an increasing trend with growing compute. Although a token-to-parameter ratio of $\approx$15 seems to be optimal in our compute regime, the results indicate that the training dataset size should scale slightly faster than the model size.
@@ Acknowledgements
Appendix A Statement of contribution
DO developed the data generation methodology, implemented all software components, conducted the formal analysis and all experiments, curated the datasets, and created the visualizations. JF conceptualised the main research idea and designed the model architecture and scaling analysis methodology. DO wrote the original draft and JF and DO performed review and editing. JF and FH supervised the project and coordinated the research activities. FF and AZ participated on preliminary discussions. FH acquired funding and provided computing resources.
Appendix B Data generation details
B.1 Expression generation
@@ B.2 Data sampling
Appendix C Additional Results
@@
Appendix D Training Details